OpenStudy (anonymous):

^3√2x3 Like the 3 on the outside is a cubed root.

OpenStudy (eyust707):

\[\sqrt[3]{2x^3}\]

OpenStudy (eyust707):

\[= \sqrt[3]{2} * \sqrt[3]{x^3}\]

OpenStudy (eyust707):

the right side is the ony part that can be simplified.. do you know what the cube root of x^3 is?

OpenStudy (anonymous):

No :/ I've never been good at maths.

OpenStudy (eyust707):

okay so cube root os x^3 means "what times what times what equals x^3"

OpenStudy (eyust707):

of*

OpenStudy (eyust707):

"what three things do i multiply by each other to get whats under the radical"

OpenStudy (eyust707):

well lets see x^3 is just x * x * x right?